S&L Financial buys and sells securities expecting to earn profits on short-term differences in price. On December 27, 2021, SS&L Financial buys and sells securities which it classifies as available-for-sale. On December 27, 2021, S&L purchased Coca-CAdams Industries holds 57,000 shares of FedEx common stock, which is not a large enough ownership interest to allow Adams toKim Company bought 40% of the shares of Phelps, Inc., at the start of 2021. Kim paid $15 million for the shares. Forty percenPlease answer all four parts ...
Answer #1
1 Computation of pre-tax amount
Year 2021
Purchase price of bond = 745,000
Fair value on 31st Dec   = 742,500
Selling Price of bond      = 748,500
As fair value is less than the purchase price, therefore there is unrealised holdig loss
Unrealised holding loss = Purchase price - Fair value
Unrealised holding loss = 745,000-742,500
Unrealised holding loss = 2,500
Year 2022
Pretax amount included in 2022
In 2022, the sale value is higher than fair value, therefore there is unrealised holding gain
Unrealised holding Gain = Selling price - Fair value
Unrealised holding Gain = 748,500-742,500
Unrealised holding Gain = 6000
2 Year 2021
Purchase price of bond = 790,000
Fair value on 31st Dec   = 786,000
Selling Price of bond      = 795,000
As fair value is less than the purchase price, therefore there is unrealised holdig loss
Unrealised holding loss = Purchase price - Fair value
Unrealised holding loss = 790,000-786,000
Unrealised holding loss = 4,000
Year 2022
Pretax amount included in 2022
In 2022, the sale value is higher than fair value, therefore there is unrealised holding gain
Unrealised holding Gain = Selling price - Fair value
Unrealised holding Gain = 795,000-786,000
Unrealised holding Gain = 9000
3 Since Adam is not holding the enough to exercise significant influence, therefore appropriate reporting category should be fair value through net income
Fair value through net income = No. of shares held X Market value as at 31st Dec 2022
Fair value through net income = 57000 X 101
Fair value through net income = $ 5,757,000
4 As Kim holds 40% in Phelps Inc.
Equity method is used if significant influence above 20% but less than 50% in the investee company
Amount of investment in Phelps Limited Amount in Million
Amount paid for shares of Phelps Limited                       15.00
Add : Share in Income                              -  
($ 2 Million * 40%)                         0.80
Less : Payment of cash dividend                              -  
($ 1.3 Million * 40%)                       (0.52)
                      15.28
Note : The difference of $3 million is due to land appreciation. It will go to the land account